# Observer-robust energy condition verification for warp drive spacetimes

### v1

@@ Abstract
We analyze five warp drive metrics (Alcubierre, Lentz, Van Den Broeck, Natário, Rodal) and one warp shell metric (used primarily as a numerical stress test). For the Rodal metric, the standard Eulerian-frame analysis misses violations at over $28\%$ of grid points (dominant energy condition) and over $15\%$ (weak energy condition). Even where the Eulerian frame identifies the correct violation set, observer optimization reveals that violation severity can be orders of magnitude larger (e.g. Alcubierre weak energy condition:  ${\sim}\,90{,}000\times$ at rapidity cap $\zeta_{\max}=5$, scaling as $e^{2\zeta_{\max}}$). These results demonstrate that single-frame evaluation can systematically underestimate both the spatial extent and the magnitude of energy condition violations in warp drive spacetimes. warpax is freely available at https://github.com/anindex/warpax.
@@ 1 Introduction
  • exact derivatives of the implemented (regularized) metric functions (up to floating-point roundoff) via forward-mode automatic differentiation, eliminating finite-difference truncation error (the computed curvature is exact for the numerical profile at each point, which may differ from the target analytic metric where smoothing or clamping is applied (see Section 3),
@@
We analyze five warp drive metrics and one warp shell geometry. For the Rodal metric, the Eulerian single-frame DEC analysis fails to detect violations at over 28% of all grid points; for WarpShell (a regularized implementation used primarily as a numerical stress test; see Section 7.3), missed WEC violations are below $0.1\%$ of the grid. For Alcubierre and Natário, we find no additional NEC or WEC violating grid points beyond the baseline Eulerian analysis (missed =0=0 at the sampled resolution); Lentz shows zero missed NEC/WEC points but nonzero SEC misses, and is severely under-resolved ($\ll\!1$ cell across the wall at $50^{3}$); this result should be treated as a lower bound. Van Den Broeck exhibits small but nonzero missed fractions (NEC $0.1\%$, WEC $0.4\%$ at $v_{s}=0.5$). The observer-optimized violation severity (at $\zeta_{\max}=5$) can be substantially larger in regions already identified as violating. We validate against WarpFactory’s published Eulerian results and quantify the additional violations discovered by observer optimization where they exist.
@@
The paper is organized as follows. Section 2 presents the mathematical methods. Section 3 describes the warpax toolkit. Section 4 presents energy condition results. Section 5 presents geodesic results. Section 6 analyses kinematic scalars. Section 7 validates numerical convergence. Section 8 discusses implications, and Section 9 concludes. Appendices provide supplementary material: transition-function ablation (Appendix A), Rodal DEC ablation (Appendix B), wall-resolution evidence (Appendix C), and full metric definitions (Appendix D).
@@ 2.2 Hawking–Ellis classification
  • Type III: A triple null eigenvector (pure radiation with no rest frame). Rare in practice.
@@
warpax implements a branchless classification algorithm suitable for GPU execution via jax.vmap. The classifier avoids Python-level conditionals by computing a continuous score for each type and selecting via branchless masking. The mixed tensor $T^{a}{}_{b}=g^{ac}T_{cb}$ is generally not symmetric in the Euclidean sense, so the eigendecomposition uses a general (non-symmetric) solver; eigenvalues are classified as real when $|\text{Im}\,\lambda_{i}|<\epsilon\,s$ (default $\epsilon=10^{-10}$, scale $s=\max(|\mathrm{Re}\,\lambda_{j}|,1)$). Eigenvector causal character and eigenvalue degeneracy use the same tolerance with the same relative scale factor. A second, relative tolerance tier catches split degenerate eigenvalue pairs at large $\|T^{a}{}_{b}\|$ (Section 8). Across all grids in this study, $\max|\text{Im}\,\lambda_{i}|/s<10^{-12}$ typically at Type I-classified points, confirming well-behaved eigenstructure (non-Type I points, including transition-zone Type IV, may have larger imaginary parts; see Table 8). Near-vacuum points ($\max|\text{Re}\,\lambda_{i}|<\epsilon$) are detected and assigned Type I with $\rho=p_{i}=0$ before the eigendecomposition, avoiding numerically meaningless eigenvectors. At exactly vacuum points ($T_{ab}=0$, e.g. Minkowski or Schwarzschild exterior), all eigenvalues are degenerate at zero and the timelike eigenvector is not unique; the classifier correctly labels these as Type I with $\rho=p_{i}=0$. For Type I, all energy conditions reduce to algebraic inequalities in the eigenvalues $\rho,p_{1},p_{2},p_{3}$:
@@
These inequalities are exact: for Type I stress-energy, a point satisfies (or violates) an energy condition if and only if the corresponding eigenvalue inequality holds, regardless of which observer is chosen. We therefore distinguish three diagnostic quantities:
@@
Throughout this paper, “robust margin” refers to the merged quantity (hereafter hybrid margin): algebraic slack at Type I points and $\zeta_{\max}$-capped extremum at non-Type I points. For Type I points, violation detection and the reported margins use the algebraic slack (item 2), which is cap-independent and authoritative. The observer optimizer is run additionally to provide the explicit worst-case observer parameters $(\zeta^{*},\theta^{*},\phi^{*})$; its $\zeta_{\max}$-capped contraction minimum (item 3) is reported separately in sensitivity analyses (Table 9) but is not used for violation detection. The algebraic slack and capped extremum are in general different numerical values: the former represents the global eigenvalue-based quantity, while the latter depends on $\zeta_{\max}$. These two quantities are conceptually distinct and should not be compared directly; the algebraic slack measures distance to the violation boundary, while the capped extremum quantifies the worst-case energy density within the tested observer range. The algebraic slack (e.g. $\min_{i}(\rho+p_{i})$ for NEC) certifies satisfaction or violation but is not, in general, equal to the minimum observer-contracted energy density $\min_{u}T_{ab}\,u^{a}u^{b}$; it measures proximity to the eigenvalue inequality boundary rather than a physical extremum over the observer manifold. For the metrics and rapidity caps used here ($\zeta_{\max}=5$) we empirically observe sign agreement between the algebraic Type I truth and the $\zeta_{\max}$-capped optimizer at all grid points (i.e. no violation is “hidden” above the cap). This is not guaranteed in general: a Type I tensor with $\rho+p_{i}<0$ but $|\rho+p_{i}|\ll|\rho|$ could require an arbitrarily large boost to flip the sign of the measured energy density, so a finite $\zeta_{\max}$ might in principle miss it. At non-Type-I grid points (up to $\sim\!3$% for the warp metrics and <1<\!1% for WarpShell; see Table 8), the algebraic inequalities do not apply and the optimizer is the sole method; this pathway is validated via a null-dust benchmark (Section 8).
@@ 2.4 Observer optimization
where $m_{\text{flux}}=-g_{ab}\,j^{a}j^{b}$ (positive when the energy flux $j^{a}=-T^{a}{}_{b}\,u^{b}$ is causal), $m_{\text{future}}=-j_{a}n^{a}$ (positive when $j^{a}$ is future-directed, with $n^{a}$ the future-pointing unit normal), and $m_{\text{WEC}}=T_{ab}\,u^{a}u^{b}$ (positive when the energy density is non-negative). A negative value of any component indicates a DEC violation for that observer. The three-term minimum in equation (17) is piecewise smooth but not $C^{1}$ at branch-switching boundaries; gradients exist almost everywhere, and the multi-start BFGS optimizer handles this reliably in practice. A log-sum-exp soft-minimum approximation could be employed if smoother gradients are needed. Because $m_{\text{WEC}}$ has dimensions of energy density while $m_{\text{flux}}$ has dimensions of (energy density)2, we do not combine them into a single margin for severity purposes. For violation detection, the DEC is violated whenever $m_{\text{DEC}}^{*}<0$ (equation 17); the sign logic is correct regardless of units. For Type I stress-energy with $\rho\geq 0$ and $\rho\geq|p_{i}|$, the algebraic condition (equation 13) guarantees that $j^{a}$ is causal and future-directed for every future-directed timelike $u^{a}$, making the three-term check redundant; the algebraic slack $\min_{i}(\rho-|p_{i}|)$ then serves as the primary DEC result with homogeneous dimensions (energy density). At non-Type I points, the optimizer evaluates all three terms of equation (17) simultaneously (Section 8). For the majority of grid points, which classify as Type I, the algebraic slack is the primary DEC result; at non-Type I points the capped extremum is used instead.
@@
Multi-start optimization is implemented via jax.vmap over $N_{\text{starts}}$ initial conditions. The first start is always the Eulerian observer ($\zeta=0$), followed by six axis-aligned spatial directions; the remaining $N_{\text{starts}}-7$ starts are drawn from a standard normal distribution. This deterministic seeding ensures the physically motivated baseline is always included. The entire multi-start pool is evaluated in parallel, which parallelises trivially on GPU hardware with no additional implementation effort.
@@ 3 The warpax toolkit
    For Type I: use algebraic slacks for violation detection; report capped extrema as $\zeta_{\max}$-dependent severity diagnostics. For non-Type-I: use capped extrema. Compute per-condition summaries.
The Eulerian-only baseline evaluates the NEC for six null vectors $k^{a}=n^{a}\pm e^{a}_{(i)}$, $i=1,2,3$, built from the Eulerian orthonormal triad $\{e^{a}_{(i)}\}$ (a restricted-direction baseline; the full NEC requires optimization over all null directions), and evaluates WEC, SEC, and DEC using the single Eulerian (ADM normal) observer. Since $m^{*}(x)=\min_{\text{observers}}\,m(x,u)\leq m(x,u_{0})$ for any fixed observer $u_{0}$, the baseline margin is an upper bound on the worst-case margin at each point, and the set of baseline-detected violations is a lower bound on the true violation extent. This inequality is enforced numerically as a consistency check in the test suite.
@@ 4 Results: Energy conditions
For each grid point, we compute: (i) the Eulerian baseline (ADM normal observer for WEC/SEC/DEC; six axis-aligned null directions for NEC), and (ii) the observer-optimized margins via multi-start optimization (Section 2.4) with $N_{\text{starts}}=8$. The Hawking–Ellis classifier identifies the vast majority of grid points as Type I (four real eigenvalues with a timelike eigenvector): Rodal achieves 100%, Lentz 99.9%, WarpShell 99.6%, VdB 98.4%, Alcubierre 98%, and Natário $\sim\!97.5$%. The minimum across all tested metrics is Natário at 97.5% (Table 8). At Type I points the algebraic eigenvalue inequalities (equations 1013) apply; at non-Type I points the optimizer provides capped extrema as the sole diagnostic.
@@ 4.1 Individual metric analysis
Figure 1: NEC evaluation for the Alcubierre metric ($50^{3}$ grid, $v_{s}=0.5$; see Table 1). Left: Eulerian margin (positive = satisfied, negative = violated). Center: Observer-optimized margin. Right: Points where the Eulerian analysis misses violations (red = missed violation). The violation sets coincide (0 missed points), but the robust margin is slightly more negative in the violating region.
Figure 2: WEC evaluation for the Alcubierre metric ($50^{3}$ grid, $v_{s}=0.5$; see Table 1). Same layout as Figure 1. The violation sets coincide (0 missed points), but the observer-optimized WEC margin at $\zeta_{\max}=5$ ($\gamma\approx 74$) is much more negative than the Eulerian value; at NEC-violating points this ratio grows with the rapidity cap.
@@
The Lentz metric [17] was originally proposed as a “positive-energy” warp drive. However, Celmaster and Rubin [21] showed that the Lentz metric violates the WEC even for Eulerian observers once the metric is properly evaluated. The diamond-pattern shift profile uses $\tanh$-based smooth transitions following the WarpFactory parameterization. Figure 3 shows the NEC comparison. Like the Alcubierre case, the observer-robust analysis finds no additional NEC violations beyond the Eulerian result (Table 3): the six-direction NEC baseline detects the full NEC-violating region at the tested parameters. The NEC violations detected by both analyses are concentrated near the bubble boundary, consistent with the findings of Celmaster and Rubin. Because the Lentz wall spans only $\sim\!0.02$ grid cells at $50^{3}$ (Table 2), the wall is effectively unresolved and any quantitative Lentz results should be regarded as qualitative indicators rather than converged values.
Figure 3: NEC evaluation for the Lentz metric ($50^{3}$ grid, $v_{s}=0.5$; see Table 1). NEC violations are concentrated near the bubble boundary. The violation sets coincide between Eulerian and robust analyses (0 missed points).
@@
The WarpShell metric models a constant-velocity physical warp drive following the Bobrick–Martire framework [22] with the Fell–Heisenberg mass parameterization [23], using WarpFactory’s “Bobrick–Martire Modified Time” simplification. It consists of a flat interior ($r<R_{1}$), a curved Schwarzschild-like shell ($R_{1}<r<R_{2}$), and an asymptotically flat exterior ($r>R_{2}$), connected by $C^{2}$-smooth quintic Hermite transition functions (Appendix A). The WarpShell metric uses a shell-like mass distribution (Schwarzschild radius parameter $r_{\mathrm{S}}=5$; smooth_width =0.06=0.06) with an inner approximately flat region, an outer asymptotically flat region (lapse and shift are exactly clamped; the spatial metric retains residual $\tanh$ tails), and a smooth transition layer. The reported WarpShell violations are properties of the regularized numerical implementation, not necessarily of an idealized thin-shell limit; the enormous NEC margin magnitudes ($\mathcal{O}(10^{30})$$\mathcal{O}(10^{35})$) reflect the regularization curvature scale and do not converge with resolution (see Section 8). The spatial metric contains a conformal factor in the shell region: $\gamma_{ij}=\delta_{ij}+(\gamma_{rr}-1)\,\hat{r}_{i}\hat{r}_{j}$ where $\gamma_{rr}$ encodes the shell mass profile. Figure 4 shows the NEC comparison. The WarpShell NEC missed fraction is negligible at $v_{s}=0.5$ (Table 3), and the WEC missed fraction is similarly small (WEC missed $0.1\%$ at $v_{s}=0.5$; DEC missed is negligible at $0.0\%$). Missed WEC violations concentrate at moderate radii ($r\approx 3.5$$4.0$), outside the shell region. The NEC and DEC violations are concentrated at the shell transition boundaries and are likely transition-zone artifacts; the WEC violations at $r\approx 3.5$$4.0$ arise from the residual $\tanh$-tail curvature and represent genuine features of the regularized geometry (see Section 8). For the <1<\!1% non-Type I WarpShell points, the three-term DEC diagnostic (equation 17) is evaluated directly, checking flux causality, future-directedness, and non-negative energy density simultaneously (Section 8).
Figure 4: NEC evaluation for the WarpShell metric (regularized implementation; $50^{3}$ grid, $v_{s}=0.5$, $R_{1}=0.5$, $R_{2}=1.0$; see Table 1). NEC and WEC missed violations are both below 0.1% at $v_{s}=0.5$ (Table 3).

### v2

@@ Abstract
We analyze five warp drive metrics (Alcubierre, Lentz, Van Den Broeck, Natário, Rodal) and one warp shell metric (used primarily as a numerical stress test). For the Rodal metric, the standard Eulerian-frame analysis misses violations at over $28\%$ of grid points (dominant energy condition) and over $15\%$ (weak energy condition). Even where the Eulerian frame identifies the correct violation set, observer optimization reveals that violation severity can be orders of magnitude larger (e.g. Alcubierre weak energy condition:  ${\sim}\,90{,}000\times$ at rapidity cap $\zeta_{\max}=5$, scaling as $e^{2\zeta_{\max}}$; the rapidity cap is an analysis hyperparameter, not a physically preferred bound). These results demonstrate that single-frame evaluation can systematically underestimate both the spatial extent and the magnitude of energy condition violations in warp drive spacetimes. warpax is freely available at https://github.com/anindex/warpax.
@@ 1 Introduction
  • exact derivatives of the implemented (regularized) metric functions (up to floating-point roundoff) via forward-mode automatic differentiation, eliminating finite-difference truncation error. Three distinct error sources should be distinguished: (a) pointwise curvature is exact for the implemented numerical profile (autodiff); (b) spatial sampling error depends on grid resolution and can miss features between grid points (Section 7); (c) model error arises where smoothing, clamping, or regularization causes the implemented profile to differ from the target analytic metric (see Section 3). Autodiff eliminates source (a) but not (b) or (c),
@@
We analyze five warp drive metrics and one warp shell geometry. For the Rodal metric (we use a lab-frame standardization $F=1-f_{\rm paper}$, $G=1-g_{\rm paper}$; see Appendix D), the Eulerian single-frame DEC analysis fails to detect violations at over 28% of all grid points; for WarpShell (a regularized implementation used primarily as a numerical stress test; see Section 7.3), missed WEC violations are below $0.1\%$ of the grid. For Alcubierre and Natário, we find no additional NEC or WEC violating grid points beyond the baseline Eulerian analysis (missed =0=0 at the sampled resolution); Lentz shows zero missed NEC/WEC points but nonzero SEC misses, and is severely under-resolved ($\ll\!1$ cell across the wall at $50^{3}$); this result should be treated as a lower bound. Van Den Broeck exhibits small but nonzero missed fractions (NEC $0.1\%$, WEC $0.4\%$ at $v_{s}=0.5$). The $\zeta_{\max}$-capped violation severity (at $\zeta_{\max}=5$) can be substantially larger in regions already identified as violating. We validate against WarpFactory’s published Eulerian results and quantify the additional violations discovered by observer optimization where they exist.
@@
The paper is organized as follows. Section 2 presents the mathematical methods. Section 3 describes the warpax toolkit. Section 4 presents energy condition results. Section 5 presents geodesic results. Section 6 analyzes kinematic scalars. Section 7 validates numerical convergence. Section 8 discusses implications, and Section 9 concludes. Appendices provide supplementary material: transition-function ablation (Appendix A), Rodal DEC ablation (Appendix B), wall-resolution evidence (Appendix C), and full metric definitions (Appendix D).
@@ 2.2 Hawking–Ellis classification
  • Type III: A triple null eigenvector (non-generic nilpotent stress–energy). Rare in practice.
@@
warpax implements a branchless classification algorithm suitable for GPU execution via jax.vmap. The classifier avoids Python-level conditionals by computing a continuous score for each type and selecting via branchless masking. The mixed tensor $T^{a}{}_{b}=g^{ac}T_{cb}$ is generally not symmetric in the Euclidean sense, so the eigendecomposition uses a general (non-symmetric) solver; eigenvalues are classified as real when $|\text{Im}\,\lambda_{i}|<\epsilon\,s$ (default $\epsilon=10^{-10}$, scale $s=\max(|\mathrm{Re}\,\lambda_{j}|,1)$). Eigenvector causal character and eigenvalue degeneracy use the same tolerance with the same relative scale factor. A second, relative tolerance tier catches split degenerate eigenvalue pairs at large $\|T^{a}{}_{b}\|$ (Section 8). Across all grids in this study, $\max|\text{Im}\,\lambda_{i}|/s<10^{-12}$ typically at Type I-classified points, confirming well-behaved eigenstructure (non-Type I points, including transition-zone Type IV, may have larger imaginary parts; see Table 8). Near-vacuum points ($\max|\text{Re}\,\lambda_{i}|<\epsilon$) are detected after the eigendecomposition and assigned Type I with $\rho=p_{i}=0$, bypassing the eigenvector-based causal-character classification whose results are numerically meaningless at near-zero eigenvalue scale. At exactly vacuum points ($T_{ab}=0$, e.g. Minkowski or Schwarzschild exterior), all eigenvalues are degenerate at zero and the timelike eigenvector is not unique; the classifier correctly labels these as Type I with $\rho=p_{i}=0$. For Type I, all energy conditions reduce to algebraic inequalities in the eigenvalues $\rho,p_{1},p_{2},p_{3}$:
@@
These inequalities are exact: for Type I stress-energy, a point satisfies (or violates) an energy condition if and only if the corresponding eigenvalue inequality holds, regardless of which observer is chosen. In floating-point arithmetic, the Type I/II/IV classification decision is tolerance-based (Section 8); all “exact” statements are in the sense of exact algebra applied to the numerically classified type, and are only as good as the eigenvalue accuracy, the complex/degeneracy tolerances, and the timelike eigenvector test. We therefore distinguish three diagnostic quantities:
@@
Throughout this paper, “robust margin” refers to the merged quantity (hereafter hybrid margin): algebraic slack at Type I points and $\zeta_{\max}$-capped extremum at non-Type I points. For Type I points, violation detection and the reported margins use the algebraic slack (item 2), which is cap-independent and authoritative. The observer optimizer is run additionally to provide the explicit worst-case observer parameters $(\zeta^{*},\theta^{*},\phi^{*})$; its $\zeta_{\max}$-capped contraction minimum (item 3) is reported separately in sensitivity analyses (Table 9) but is not used for violation detection. The algebraic slack and capped extremum are in general different numerical values: the former represents the global eigenvalue-based quantity, while the latter depends on $\zeta_{\max}$. These two quantities are conceptually distinct and should not be compared directly; the algebraic slack measures distance to the violation boundary, while the capped extremum quantifies the worst-case energy density within the tested observer range. The algebraic slack (e.g. $\min_{i}(\rho+p_{i})$ for NEC) certifies satisfaction or violation but is not, in general, equal to the minimum observer-contracted energy density $\min_{u}T_{ab}\,u^{a}u^{b}$; it measures proximity to the eigenvalue inequality boundary rather than a physical extremum over the observer manifold. For the metrics and rapidity caps used here ($\zeta_{\max}=5$) we empirically observe sign agreement between the algebraic Type I truth and the $\zeta_{\max}$-capped optimizer at all grid points (i.e. no violation is “hidden” above the cap). This is not guaranteed in general: a Type I tensor with $\rho+p_{i}<0$ but $|\rho+p_{i}|\ll|\rho|$ could require an arbitrarily large boost to flip the sign of the measured energy density, so a finite $\zeta_{\max}$ might in principle miss it. At non-Type I grid points (up to $\sim\!3$% for the warp metrics and <1<\!1% for WarpShell; see Table 8), the algebraic inequalities do not apply and the optimizer is the sole method; this pathway is validated via a null-dust benchmark (Section 8).
@@ 2.4 Observer optimization
where $m_{\text{flux}}=-g_{ab}\,j^{a}j^{b}$ (positive when the energy flux $j^{a}=-T^{a}{}_{b}\,u^{b}$ is causal), $m_{\text{future}}=-j_{a}n^{a}$ (positive when $j^{a}$ is future-directed, with $n^{a}$ the future-pointing unit normal), and $m_{\text{WEC}}=T_{ab}\,u^{a}u^{b}$ (positive when the energy density is non-negative). A negative value of any component indicates a DEC violation for that observer. The three-term minimum in equation (17) is piecewise smooth but not $C^{1}$ at branch-switching boundaries; gradients exist almost everywhere, and the multi-start BFGS optimizer handles this reliably in practice. A log-sum-exp soft-minimum approximation could be employed if smoother gradients are needed. An alternative design would run three separate optimizations (one per DEC sub-condition) and take the minimum of the minima; this would avoid the unit-mixing landscape and improve interpretability of the “worst-case observer” at some additional computational cost. Because $m_{\text{WEC}}$ has dimensions of energy density while $m_{\text{flux}}$ has dimensions of (energy density)2, we do not combine them into a single margin for severity purposes. For violation detection, the DEC is violated whenever $m_{\text{DEC}}^{*}<0$ (equation 17); the sign logic is correct regardless of units. For Type I stress-energy with $\rho\geq 0$ and $\rho\geq|p_{i}|$, the algebraic condition (equation 13) guarantees that $j^{a}$ is causal and future-directed for every future-directed timelike $u^{a}$, making the three-term check redundant; the algebraic slack $\min_{i}(\rho-|p_{i}|)$ then serves as the primary DEC result with homogeneous dimensions (energy density). At non-Type I points, the optimizer evaluates all three terms of equation (17) simultaneously (Section 8). For the majority of grid points, which classify as Type I, the algebraic slack is the primary DEC result; at non-Type I points the capped extremum is used instead.
@@
Multi-start optimization is implemented via jax.vmap over $N_{\text{starts}}$ initial conditions. The first start is always the Eulerian observer ($\zeta=0$), followed by six axis-aligned spatial directions; the remaining $N_{\text{starts}}-7$ starts are drawn from a standard normal distribution. This deterministic seeding ensures the physically motivated baseline is always included. The entire multi-start pool is evaluated in parallel, which parallelizes trivially on GPU hardware with no additional implementation effort.
@@ 3 The warpax toolkit
    For Type I: use algebraic slacks for violation detection; report capped extrema as $\zeta_{\max}$-dependent severity diagnostics. For non-Type I: use capped extrema. Compute per-condition summaries.
The Eulerian-only baseline evaluates the NEC for six null vectors $k^{a}=n^{a}\pm e^{a}_{(i)}$, $i=1,2,3$, built from the Eulerian orthonormal triad $\{e^{a}_{(i)}\}$. This is a restricted-direction lower bound on the NEC violation set (six rays, not the full $S^{2}$); it can miss violations whose worst-case null direction falls between the sampled axes. The baseline evaluates WEC, SEC, and DEC using the single Eulerian (ADM normal) observer. Since $m^{*}(x)=\min_{\text{observers}}\,m(x,u)\leq m(x,u_{0})$ for any fixed observer $u_{0}$, the baseline margin is an upper bound on the worst-case margin at each point, and the set of baseline-detected violations is a lower bound on the true violation extent. This inequality is enforced numerically as a consistency check in the test suite.
@@ 4 Results: Energy conditions
For each grid point, we compute: (i) the Eulerian baseline (ADM normal observer for WEC/SEC/DEC; six axis-aligned null directions for NEC), and (ii) the hybrid margins (algebraic slack at Type I points, $\zeta_{\max}$-capped observer extremum at non-Type I) via multi-start optimization (Section 2.4) with $N_{\text{starts}}=8$. The Hawking–Ellis classifier identifies the vast majority of grid points as Type I (four real eigenvalues with a timelike eigenvector): Rodal achieves 100%, Lentz 99.9%, WarpShell 99.6%, VdB 98.4%, Alcubierre 98%, and Natário $\sim\!97.5$%. The minimum across all tested metrics is Natário at 97.5% (Table 8). At Type I points the algebraic eigenvalue inequalities (equations 1013) apply; at non-Type I points the optimizer provides capped extrema as the sole diagnostic.
@@ 4.1 Individual metric analysis
Figure 1: NEC evaluation for the Alcubierre metric ($50^{3}$ grid, $v_{s}=0.5$; see Table 1). Left: Eulerian margin (positive = satisfied, negative = violated). Center: Hybrid margin (algebraic slack at Type I points, $\zeta_{\max}$-capped extremum at non-Type I). Right: Points where the Eulerian analysis misses violations (red = missed violation). This panel serves as a sanity check: for the Alcubierre NEC, the violation sets coincide (0 missed points), confirming that the observer-robust method recovers the Eulerian result when no additional violations exist. Non-trivial missed violations are demonstrated for other metrics and conditions (Figures 34, Table 3).
Figure 2: WEC evaluation for the Alcubierre metric ($50^{3}$ grid, $v_{s}=0.5$; see Table 1). Same layout as Figure 1. The violation sets coincide (0 missed points), but the hybrid WEC margin at $\zeta_{\max}=5$ ($\gamma\approx 74$) is much more negative than the Eulerian value; at NEC-violating points this ratio grows with the rapidity cap.
@@
The Lentz metric [17] was originally proposed as a “positive-energy” warp drive. However, Celmaster and Rubin [21] showed that the Lentz metric violates the WEC even for Eulerian observers once the metric is properly evaluated. The diamond-pattern shift profile uses $\tanh$-based smooth transitions following the WarpFactory parameterization. Like the Alcubierre case, the observer-robust analysis finds no additional NEC violations beyond the Eulerian result (Table 3): the six-direction NEC baseline detects the full NEC-violating region at the tested parameters. The NEC violations detected by both analyses are concentrated near the bubble boundary, consistent with the findings of Celmaster and Rubin. Because the Lentz wall spans only $\sim\!0.02$ grid cells at $50^{3}$ (Table 2), the wall is effectively unresolved and any quantitative Lentz results should be regarded as qualitative indicators rather than converged values.
Van Den Broeck metric.
The Van Den Broeck metric modifies Alcubierre’s construction by introducing a conformal factor $B(r_{s})$ on the spatial metric, expanding the bubble’s internal volume while keeping the external surface area small. Unlike Alcubierre and Lentz, the observer-robust analysis discovers missed NEC, WEC, SEC, and DEC violations (Table 3), with SEC showing the largest conditional miss rate ($f_{\text{miss}|\text{viol}}=17.5\%$; Figure 4). The NEC comparison (Figure 3) confirms that even small miss fractions ($0.1\%$ of grid points) represent genuine additional violations that the Eulerian frame fails to detect.
Figure 3: NEC evaluation for the Van Den Broeck metric ($50^{3}$ grid, $v_{s}=0.5$; see Table 1). Left: Eulerian margin. Center: Hybrid margin (algebraic slack at Type I, capped extremum at non-Type I). Right: Missed violations (red). The observer-robust analysis finds $0.1\%$ of grid points where the Eulerian frame misses NEC violations (conditional miss rate $f_{\text{miss}|\text{viol}}=1.6\%$; Table 3).
@@
The WarpShell metric models a constant-velocity physical warp drive following the Bobrick–Martire framework [22] with the Fell–Heisenberg mass parameterization [23], using WarpFactory’s “Bobrick–Martire Modified Time” simplification. It consists of a flat interior ($r<R_{1}$), a curved Schwarzschild-like shell ($R_{1}<r<R_{2}$), and an asymptotically flat exterior ($r>R_{2}$), connected by $C^{2}$-smooth quintic Hermite transition functions (Appendix A). The WarpShell metric uses a shell-like mass distribution (Schwarzschild radius parameter $r_{\mathrm{S}}=5$; smooth_width =0.06=0.06) with an inner approximately flat region, an outer asymptotically flat region (lapse and shift are exactly clamped; the spatial metric retains residual $\tanh$ tails), and a smooth transition layer. Because $r_{\mathrm{S}}>R_{2}>R_{1}$, the Schwarzschild expressions require clamping regularization to remain real-valued (see Appendix D, “WarpShell regularization”). The reported WarpShell violations are properties of the regularized numerical implementation, not necessarily of an idealized thin-shell limit; the enormous NEC margin magnitudes ($\mathcal{O}(10^{30})$$\mathcal{O}(10^{35})$) reflect the regularization curvature scale and do not converge with resolution (see Section 8). The spatial metric contains a conformal factor in the shell region: $\gamma_{ij}=\delta_{ij}+(\gamma_{rr}-1)\,\hat{r}_{i}\hat{r}_{j}$ where $\gamma_{rr}$ encodes the shell mass profile. The WarpShell NEC missed fraction is negligible at $v_{s}=0.5$ (Table 3), and the WEC missed fraction is similarly small (WEC missed $0.1\%$ at $v_{s}=0.5$; DEC missed is negligible at $0.0\%$). Missed WEC violations concentrate at moderate radii ($r\approx 3.5$$4.0$), outside the shell region. The NEC and DEC violations are concentrated at the shell transition boundaries and are likely transition-zone artifacts; the WEC violations at $r\approx 3.5$$4.0$ arise from the residual $\tanh$-tail curvature and represent genuine features of the regularized geometry (see Section 8). For the <1<\!1% non-Type I WarpShell points, the three-term DEC diagnostic (equation 17) is evaluated directly, checking flux causality, future-directedness, and non-negative energy density simultaneously (Section 8).
Figure 4: SEC evaluation for the Van Den Broeck metric ($50^{3}$ grid, $v_{s}=0.5$; see Table 1). Left: Eulerian margin. Center: Hybrid margin (algebraic slack at Type I, capped extremum at non-Type I). Right: Missed violations (red). SEC exhibits the largest conditional miss rate among all conditions for Van Den Broeck ($f_{\text{miss}|\text{viol}}=17.5\%$, $1.2\%$ of grid points missed; Table 3), demonstrating that SEC violations are genuinely independent of the NEC/WEC/DEC hierarchy.